# RemoteReasoner: Towards Unifying Geospatial Reasoning Workflow

### v2

@@ 4.2 Implementations Details
We employ GRPO Shao et al. (2024) to fine-tune the Qwen2.5-VL-7B-Instruct Bai et al. (2025) model using LoRA Hu et al. (2022) (rank=8, alpha=16, targeting all linear layers). Training is conducted for 24 epochs on 8 × NVIDIA L20 GPUs with a global batch size of 512 (achieved via per-device batch size 8 and gradient accumulation steps 8), using the AdamW Loshchilov and Hutter (2017) optimizer with a learning rate of 1e-6, bf16 precision, and DeepSpeed Rasley et al. (2020) Zero3 for memory efficiency.
@@ 4.3 Main Results
4.3.1 Geospatial Region Reasoning
This section validates RemoteReasoner’s performance on region-level reasoning tasks, with results presented in Tab. 2. Compared to the remote sensing multi-modal foundation model GeoChat Kuckreja et al. (2024), RemoteReasoner demonstrates at least a 56.41% improvement. This significant gain stems from GeoChat’s weak reasoning capabilities, which fail to interpret ambiguous queries. Similarly, RemoteReasoner achieves over 22.39% improvement against general-purpose MLLMs, e.g., Qwen-2.5VL-7B Bai et al. (2025) and DeepSeek-VL2-tiny (3B) Guo et al. (2025). Although these MLLMs possess basic reasoning abilities, their insufficient domain-specific knowledge in remote sensing leads to suboptimal performance on geospatial region-level reasoning tasks.
4.3.2 Geospatial Pixel Reasoning
We evaluate RemoteReasoner’s performance on pixel-level reasoning tasks using the EarthReason dataset. We compare against state-of-the-art general-domain methods (LISA Lai et al. (2024), PixelLM Ren et al. (2024), PSALM Zhang et al. (2024)) and the remote sensing-domain SegEarth-R1 Li et al. (2025), with results summarized in Tab. 1. Experimental results demonstrate that despite not directly utilizing pixel-level labels during training, our approach remarkably surpasses SFT-based methods that leverage both pixel annotations and reasoning trajectories. For instance, RemoteReasoner achieves 10.41% higher $cIoU$ than LISA on the validation set. Compared to the best-performing SegEarth-R1, we observe a 3.67% improvement. The primary reason lies in the superior robustness of reinforcement learning (RL) frameworks compared to supervised fine-tuning (SFT) paradigms. Even without direct exposure to pixel-level annotations, our model effectively utilizes bounding boxes to prompt the Segment Anything Model (SAM), achieving precise segmentation results through this indirect yet efficient interaction mechanism.
4.3.3 Geospatial Contour Reasoning
This section validates the contour extraction performance of RemoteReasoner. We select EPOC Chen et al. (2024a) as a comparative baseline. Since EPOC lacks reasoning capabilities, we crop the bounding boxes (bbox) generated by RemoteReasoner and input them into EPOC to obtain contours, as shown in Fig. 3. Since EPOC is class-agnostic, it extracts contours for all objects within the region (e.g., first and third rows). In contrast, RemoteReasoner’s contours are derived from segmentation masks, resulting in tighter alignment with target boundaries. Furthermore, under low-resolution conditions (second row), EPOC produces blurry or disconnected contours, whereas our method maintains sharp and closed boundaries.
@@ 4.3.3 Geospatial Contour Reasoning
Figure 3: Geospatial Contour Reasoning Results. We select EPOC Chen et al. (2024a) for comparison.
@@
MGVLF Zhan et al. (2023) 76.78 68.04
Falcon Yao et al. (2025a) 56.87 -
SkyEyeGPT Zhan et al. (2024) 70.50 -
LHRS-Bot Muhtar et al. (2024) 73.45 -
@@
MiniGPT-V2 Zhu et al. (2023) 46.64 -
Qwen-VL-Chat Hui et al. (2024) 44.76 -
@@
LISA Lai et al. (2024) - 26.78
PixelLM Ren et al. (2024) - 31.65
@@
Figure 5: Quantitative Results. Given an image and its corresponding implicit query, RemoteReasoner autonomously identifies the user-intended target category through reasoning and accurately executes visual-centric tasks across three granularities.
@@ 4.4 Further Analysis
4.4.1 Can RemoteReasoner Do Other Tasks?
We evaluate RemoteReasoner on Visual Grounding and Referring Expression Segmentation tasks, with results summarized in Tab. 4. Comparative experiments are conducted against state-of-the-art LLM-based approaches, including general-domain methods (MiniGPT-V2 Zhu et al. (2023), LISA Lai et al. (2024), PixelLM Ren et al. (2024)) and remote sensing-specific frameworks (Falcon Yao et al. (2025a), LHRS-Bot Muhtar et al. (2024), GeoGround Zhou et al. (2024)). Notably, all remote sensing methods were trained on task-specific data, while RemoteReasoner’s zero-shot performance achieves 80% of their supervised results. Under identical zero-shot settings, our framework achieves gains of 4.12% on Visual Grounding and 19.32% on Referring Expression Segmentation tasks.
@@ 4.4.1 Can RemoteReasoner Do Other Tasks?
4.4.2 Can RemoteReasoner Recognize Other Category?
Given that the EarthReason Li et al. (2025) dataset contains only 28 categories, which is significantly insufficient to cover common remote sensing targets. Therefore, we further investigate RemoteReasoner’s recognition capability for unseen categories. We manually select 100 images spanning 10 novel categories from RemoteSAM-270K Yao et al. (2025b) that are absent in EarthReason for evaluation. As represented in Tab. 5, RemoteReasoner demonstrates robust recognition of out-of-distribution categories with consistent accuracy. It demonstrates that our RL strategy avoids overfitting to category distributions in the training data. However, the model still exhibits insufficient capabilities in detecting small objects (e.g., , cars), indicating a direction for future research.
4.4.3 Qualitative Analysis
@@ 4.5 Ablation Studies
4.5.1 Different Pixel-level Output Formats
Recently, RemoteSAM Yao et al. (2025b) validates that segmentation, as a fundamental visual output unit, enables bottom-up transformation to other vision tasks. Therefore, if LLMs could directly generate masks in textual representations, the architectural design holds promise for greater unification. To explore this hypothesis, we conducted experiments using such formulations, as summarized in Tab. 6. However, textual outputs prove ill-suited for this predictive task, with both SFT and RL strategies exhibiting significant performance degradation compared to visual-based approaches. In contrast, directly outputting bounding boxes and then generating masks through SAM significantly outperforms textual mask representations (22.01 vs. 60.25). When incorporating RL strategies, the accuracy further improves from 60.25 to 69.13. Based on these findings, we adopt the Mask2Contour approach in our workflow design.
4.5.2 Data Scaling Experiment
@@
4.5.3 Inference Cost Ablation
@@ 5 Conclusion & Limitations
Conclusion. We present RemoteReasoner, a novel workflow advancing autonomous geospatial reasoning. Overcoming limitations of SFT-based approaches, it leverages RL strategy to cultivate robust reasoning capabilities while preserving MLLM generalization. Its core innovation is a unified inference pipeline enabling efficient multi-granularity task transformation (pixel-, region-, and object-level) from a single MLLM output, eliminating redundant per-task decoders. RemoteReasoner achieves SOTA results, including >> 15% and 30% accuracy gains in region- and object-level reasoning, and demonstrates superior generalization across diverse, complex queries. It provides a flexible foundation for downstream geospatial intelligence requiring nuanced intent interpretation and multi-format outputs.
Limitations. Despite the advantages, several limitations exist for future improvement. First, model performance remains limited by current reasoning dataset scale. Progress will require more comprehensive datasets. Second, the performance on small objects needs enhancement via specialized data and architectural refinements. Additionally, exploring more tasks could advance unified geospatial reasoning.
Appendix of RemoteReasoner
A. More Implementation Details
The training pipeline leverages MS-SWIFT framework with Group Relative Preference Optimization (GRPO). Experiments use 8 $\times$ NVIDIA L20 GPUs with DeepSpeed Zero-3 for memory optimization. We initialize weights from Qwen2.5-VL-7B-Instruct and apply LoRA adaptation (rank 8, $\alpha=16$) to all linear layers. The training configuration employs a global batch size of 512 (8 per-device batches $\times$ 8 gradient accumulation steps $\times$ 8 GPUs), bfloat16 precision, and flash attention for computational efficiency. We train for 24 epochs with a learning rate of $1\times 10^{-6}$, linear warmup (5% of steps), and gradient checkpointing. Input sequences are capped at 4096 tokens with image resolution limited to 802816 pixels. The optimizer uses AdamW with weight decay disabled. For RLHF alignment, we utilize a temperature of 0.9 during generation.
B. More Experiments
B.1 Balance Factors Sensitive Analysis
Table 7: Balance Factors Sensitive Analysis of Accuracy Rewards.
gIoUgIoU 0.3 0.5 0.7 1
0.3 47.24 46.73 46.22 48.01
0.5 45.21 46.85 48.03 49.26
0.7 47.07 46.90 49.88 49.05
1 47.78 49.67 48.95 50.11
Before training, to obtain an optimal set of balance factors for the reward function, we sampled a small portion of the dataset and trained the model using different weighting schemes for the accuracy rewards. Tab. 7 presents a sensitivity analysis of the reward function balance factors, each ranging from 0.3 to 1.0. The values represent model performance under each configuration on the validation set. Among all tested combinations, the setting where both balance factors are set to 1.0 achieves the best performance, reaching 50.11%. Compared to lower-weighted settings, the (1.0, 1.0) configuration provides stronger and more consistent supervision, leading to more effective optimization. Therefore, we adopt this configuration in our final model training.
B.2 Point Prompt Effectiveness Ablations
Table 8: Effectiveness of different prompts to SAM. We directly calculate bbox’s center point without letting LLM output additional points.
Bbox Point cIoUcIoU gIoUgIoU
$\checkmark$ 43.09 42.41
$\checkmark$ $\checkmark$ 69.13 70.96
To evaluate the effectiveness of different prompt types in guiding the SAM model, we conducted a controlled experiment using bounding box (bbox) and center point prompts. Specifically, we compare two settings: using only bbox prompts, and using both bbox and point prompts, where the point is simply computed as the center of the bbox—without relying on additional points generated by the LLM. As shown in the Tab. 8, using only the bbox yields moderate performance with 43.09 $cIoU$ and 42.41 $gIoU$. However, when the center point is additionally provided as a prompt, performance improves significantly to 69.13 $cIoU$ and 70.96 $gIoU$. This substantial gain highlights the complementary role of point prompts in guiding the segmentation process. The result suggests that even simple, derived point inputs (like bbox centers) can substantially enhance SAM’s segmentation accuracy when used in conjunction with bounding boxes.
C. More Qualitative Results
In this section, we present additional visualizations for multi-granularity reasoning tasks. Results for region-level and pixel-level reasoning are illustrated in Fig. 7, while contour reasoning outcomes appear in Fig. 8. Results incorporating the reasoning process are depicted in Fig. 9 and  10. Additionally, we provide image-level task results—including Visual Question Answering (VQA) and image captioning—as summarized in Fig. 11 and  12.
Figure 7: Region- (detection) and Pixel-level (segmentation) Results.
Figure 8: Object-level (contour extraction) Results.
Figure 9: More Geospatial Reasoning Results (part 1).
Figure 10: More Geospatial Reasoning Results (part 2).
Figure 11: More Qualitative Results of Visual Question Answering.
Figure 12: More Qualitative Results of Image Captioning.
D Failure Case Analysis
Fig. 13 presents several illustrative failure cases of RemoteReasoner. First row: The query requests an entity that circumvents the urban core (expected: an overpass). However, the model incorrectly classifies the entire intersection as fulfilling this condition, consequently outputting its complete structure. Second row: The query requires locating the grandstand. While the model outputs the entire baseball stadium, the reasoning process leading to this output is logically sound. It is noteworthy that the ground truth annotation for this sample is also incorrect, as it fails to label the grandstand location. Third row: The model mistakenly incorporates the shadow region as part of the target object. These failure cases highlight significant areas for improvement in our method.
Figure 13: Failure Case Qualitative Results.
E. Prompts Setting
We present the prompt templates designed during dataset construction in Tab. 9.
Table 9: Prompt setting used for GRPO training. The text description and assistant output shown are examples; actual inputs and outputs vary by sample.
Prompt for Bounding Box Prediction
System Message:
Task Definition: You are a remote sensing image understanding and reasoning assistant. Your task is to locate the target object in the image based on the provided text description and output its bounding box coordinates. Please follow these requirements: 1. Text comprehension: Carefully analyze the text description to identify the target object. 2. Output format: If one object exists: Output the bounding box coordinates in the format: [x_min, y_min, x_max, y_max]. If no matching object is found: Output an empty list []. If multi-objects exist: Output the bounding box coordinates in the format: [[x_min1, y_min1, x_max1, y_max1], [x_min2, y_min2, x_max2, y_max2], ...]. 3. Coordinate requirements: Coordinate values must be integers representing pixel positions. Critical: You MUST output the tight bounding box around the object. 4. Output specifications: Output the thinking process in <think> </think> and final answer (a list of bbox) in <answer> </answer> tags. Task Input:
Perform the detection task for the following input:
<image>
Text description: (A natural language instruction describing the target object.)
User:
(This is an example question. Actual descriptions should be adapted to specific task needs.)
Assistant:
Reasoning Process: <think> The airport helipad is a designated area for safe helicopter landings and departures, clearly marked to ensure exclusively secure operations for aerial vehicles. </think> Bounding Box Output: <answer>[[500,164,605,246],[404,320,516,404]]</answer>

### v3

@@ 4.2 Implementations Details
We employ GRPO (shao2024deepseekmath) to fine-tune the Qwen2.5-VL-7B-Instruct (bai2025qwen2) model using LoRA (hu2022lora) (rank=8, alpha=16, targeting all linear layers). Training is conducted for 24 epochs on 8 × NVIDIA L20 GPUs with a global batch size of 512 (achieved via per-device batch size 8 and gradient accumulation steps 8), using the AdamW (loshchilov2017decoupled) optimizer with a learning rate of 1e-6, bf16 precision, and DeepSpeed (rasley2020deepspeed) Zero3 for memory efficiency.
@@ 4.3 Main Results
Geospatial Pixel Reasoning
We evaluate RemoteReasoner’s performance on pixel-level reasoning tasks using the EarthReason dataset. We compare against state-of-the-art general-domain methods (LISA (lai2024lisa), PixelLM (ren2024pixellm), PSALM (zhang2024psalm)) and the remote sensing-domain SegEarth-R1 (li2025segearth), with results summarized in Tab. 1. Experimental results demonstrate that despite not directly utilizing pixel-level labels during training, our approach remarkably surpasses SFT-based methods that leverage both pixel annotations and reasoning trajectories. For instance, RemoteReasoner achieves 10.41% higher $cIoU$ than LISA on the validation set. Compared to the best-performing SegEarth-R1, we observe a 3.67% improvement. The primary reason lies in the superior robustness of reinforcement learning (RL) frameworks compared to supervised fine-tuning (SFT) paradigms. Even without direct exposure to pixel-level annotations, our model effectively utilizes bounding boxes to prompt the Segment Anything Model (SAM), achieving precise segmentation results through this indirect yet efficient interaction mechanism.
Geospatial Region Reasoning
This section validates RemoteReasoner’s performance on region-level reasoning tasks, with results presented in Tab. 2. Compared to the remote sensing multi-modal foundation model GeoChat (kuckreja2024geochat), RemoteReasoner demonstrates at least a 54.47% improvement in $Acc@0.5$. This significant gain stems from GeoChat’s weak reasoning capabilities, which fail to interpret ambiguous queries. Similarly, RemoteReasoner achieves over 22.29% improvement against general-purpose MLLMs, e.g., Qwen-2.5VL-7B (bai2025qwen2) and DeepSeek-VL2-tiny (3B) (guo2025deepseek). Although these MLLMs possess basic reasoning abilities, their insufficient domain-specific knowledge in remote sensing leads to suboptimal performance on geospatial region-level reasoning tasks.
Geospatial Contour Reasoning
This section validates the contour extraction performance of RemoteReasoner. We select EPOC (chen2024subobject) as a comparative baseline. Since EPOC lacks reasoning capabilities, we crop the bounding boxes (bbox) generated by RemoteReasoner and input them into EPOC to obtain contours, as shown in Fig. 3. As EPOC is class-agnostic, it extracts contours for all objects within the region (e.g., Fig. 3 (a) and (b)). In contrast, RemoteReasoner’s contours are derived from segmentation masks, resulting in tighter alignment with target boundaries. Furthermore, under low-resolution conditions (Fig. 3 (c)), EPOC produces blurry or disconnected contours, whereas our method maintains sharp and closed boundaries.
@@
MGVLF 76.78 68.04
Falcon 56.87 -
SkyEyeGPT 70.50 -
LHRS-Bot 73.45 -
@@
MiniGPT-V2 46.64 -
Qwen-VL-Chat 44.76 -
@@
LISA - 26.78
PixelLM - 31.65
NEXT-Chat - 24.98
@@
Figure 5: Qualitative Results. Given an image and its corresponding implicit query, RemoteReasoner autonomously identifies the user-intended target category through reasoning and accurately executes visual-centric tasks across three granularities.
@@ 4.4 Further Analysis
Can RemoteReasoner Do Other Tasks?
We evaluate RemoteReasoner on Visual Grounding and Referring Expression Segmentation tasks, with results summarized in Tab. 4. Comparative experiments are conducted against state-of-the-art LLM-based approaches, including general-domain methods (MiniGPT-V2 (zhu2023minigpt), LISA (lai2024lisa), PixelLM (ren2024pixellm)) and remote sensing-specific frameworks (Falcon (yao2025falcon), LHRS-Bot (muhtar2024lhrs), GeoGround (zhou2024geoground)). Notably, all remote sensing methods were trained on task-specific data, while RemoteReasoner’s zero-shot performance achieves 80% of their supervised results. Under identical zero-shot settings, our framework achieves gains of 4.12% on Visual Grounding and 19.32% on Referring Expression Segmentation tasks.
@@
Can RemoteReasoner Recognize Other Category?
Given that the EarthReason (li2025segearth) dataset contains only 28 categories, which is significantly insufficient to cover common remote sensing targets. Therefore, we further investigate RemoteReasoner’s recognition capability for unseen categories. We manually select 100 images spanning 10 novel categories from RemoteSAM-270K (yao2025remotesam) that are absent in EarthReason for evaluation. As represented in Tab. 5, RemoteReasoner demonstrates robust recognition of out-of-distribution categories with consistent accuracy. It demonstrates that our RL strategy avoids overfitting to category distributions in the training data. However, the model still exhibits insufficient capabilities in detecting small objects (e.g., cars), indicating a direction for future research.
Qualitative Analysis
@@ 4.5 Ablation Studies
Different Pixel-level Output Formats
Recently, RemoteSAM (yao2025remotesam) validates that segmentation, as a fundamental visual output unit, enables bottom-up transformation to other vision tasks. Therefore, if LLMs could directly generate masks in textual representations, the architectural design holds promise for greater unification. To explore this hypothesis, we conducted experiments using such formulations, as summarized in Tab. 6. However, textual outputs prove ill-suited for this predictive task, with both SFT and RL strategies exhibiting significant performance degradation compared to visual-based approaches. In contrast, directly outputting bounding boxes and then generating masks through SAM significantly outperforms textual mask representations (22.01 vs. 60.25). When incorporating RL strategies, the accuracy further improves from 60.25 to 69.13. Based on these findings, we adopt the Mask2Contour approach in our workflow design.
Data Scaling Experiment
@@
Inference Cost Ablation
@@ 5 Conclusion
We present RemoteReasoner, a novel workflow advancing autonomous geospatial reasoning. Overcoming limitations of SFT-based approaches, it leverages RL strategy to cultivate robust reasoning capabilities while preserving MLLM generalization. Its core innovation is a unified inference pipeline enabling efficient multi-granularity task transformation (pixel-, region-, and object-level) from a single MLLM output, eliminating redundant per-task decoders. RemoteReasoner achieves SOTA results, including >> 15% and 30% accuracy gains in region- and object-level reasoning, and demonstrates superior generalization across diverse, complex queries. It provides a flexible foundation for downstream geospatial intelligence requiring nuanced intent interpretation and multi-format outputs.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 62372155 and Grant 62302149, in part by Basic Research Program of Jiangsu under Grant BK20250188, in part by the Research Funds of Jiangsu Hydraulic Research Institute under Grant 2025z065, and in part by Changzhou Science and Technology Bureau Project No. 20231313, and in part by Research and Application of the Next-Generation General-Purpose Intelligent Robot Brain (Robo-GPT).